If 150 mL of Ca(OH)2 is titrated by 30 mL of 6 M HCl, what is the concentration of the calcium hydroxide?
1
Expert's answer
2017-01-26T05:10:49-0500
Solution:
1) Write down the balanced equation:
Ca(OH)2 + 2HCl = CaCl2 +2H2O
We can see that one mole of calcium hydroxide requires 2 moles of HCl for neutralization.
2) Let’s find number of moles in titrant:
Molar concentration (molarity) shows number of moles of solute in 1000 mL of solution. So 1000 mL of our HCl solution contains 6 moles of HCl.
Than 30 mL of such solution contains 30 * 6 /1000 = 0.18 moles of HCl.
3) Find the concentration of calcium hydroxide:
If 2 moles of acid titrate 1 mole of calcium hydroxide than 0.18 moles of acid titrate 0.18 * 1 / 2 = 0.09 moles of calcium hydroxide. So 150 mL of calcium hydroxide solution contain 0.09 moles of the solute, than 1000 mL of such solution would contain 1000 * 0.09 / 150 = 0.6 moles. So we have 0.6 M solution of Ca(OH)2.

Answer:
The concentration of the calcium hydroxide is 0.6 M (mol/L)